# Generative Psycho-Lexical Approach for Constructing Value Systems in Large Language Models

### v1

@@ 1 Introduction
Personal values are broad, fundamental motivations behind individual and collective actions. They serve as guiding principles that shape perceptions, cognition, and behaviors across situations and over time. As Large Language Models (LLMs) continue to permeate and transform human society, researchers have increasingly sought to evaluate, understand, and align LLM values [92, 65, 64, 55, 44, 37, 67, 90, 94, 82, 56, 19]. Rather than relying solely on specific risk metrics, implicit preference modeling, or atomic ethical principles, framing these studies through the lens of intrinsic values [76] offers a more comprehensive, adaptable, and transparent approach [89, 9, 88].
@@
However, value systems used in existing research—primarily Schwartz’s values—are designed for humans. A psychologically informed value system specifically tailored for LLMs remains largely unexplored, obscuring a structured and holistic perspective on their values. A preliminary attempt at constructing LLM value systems [9] has several limitations that undermine its psychological grounding; we present related discussion and experimental evidence in § 2.3 and Appendix C.
@@
To address this gap, we introduce a generative psycho-lexical approach (GPLA), a novel methodology that leverages LLMs to construct a value system grounded in psychological principles. This approach is based on the psycho-lexical hypothesis, which suggests that all salient values are captured in language. GPLA adopts an agentic framework and involves five sequential steps: 1) extracting perceptions from the corpus, 2) identifying values behind perceptions, 3) filtering values, 4) non-reactive value measurements of test subjects, and 5) structuring the value system. In contrast to traditional methods that manually compile value lexicons and design questionnaires to collect self-report data [1, 16, 42, 72], GPLA automates the entire process and features two key advantages.
@@
First, GPLA supports agent-specific construction. Rather than relying on value lexicons derived from human-compiled dictionaries, it utilizes LLMs to identify values within free-form agent-generated text. This principled approach facilitates the agent- and context-specific collection of value lexicons and ensures better coverage (§ B.1). Second, GPLA enables nonreactive value measurement. Traditional methods that rely on forced-choice self-report are demonstrated to be unreliable and biased for humans [60, 8] and AI [92, 66, 21]. Instead, GPLA measures values through analyzing unstructured responses to arbitrary prompts, which mitigates response biases and enables scalable and self-adaptive value measurement [92].
@@
In applying GPLA to 33 LLMs, each steered with 21 different value-anchoring prompts [67], we develop a five-factor value system encompassing Social Responsibility, Risk-taking, Rule-following, Self-competence, and Rationality. We demonstrate the system’s reliability and explore its theoretical and practical implications. To benchmark LLM value systems, we propose three tasks based on psychometric principles [20] and AI priorities [34, 3]: Confirmatory Factor Analysis for assessing structural validity, LLM Safety Prediction for examining predictive validity, and LLM Value Alignment for evaluating representational power. These tasks confirm the validity and utility of our system in evaluating, understanding, and aligning the intrinsic values of LLMs, when compared to canonical Schwartz’s values [76].
@@ 2.1 Values, Value Measurement, and Theory-Driven Value Systems
Values are broad, fundamental beliefs or principles that guide individuals’ behaviors, cognition, and decisions [78]. They represent what people consider important in life and serve as motivational forces when evaluating actions, people, and events [76]. Unlike other psychological constructs such as personalities, values are stable, motivational, and transituational [70]. These characteristics make values uniquely powerful for understanding human behaviors [7], evaluating individual and cultural differences [74], and resolving conflicts rooted in clashing priorities [81].
@@
Values are measurable constructs [80]. Value measurement seeks a quantitative assessment of the importance individuals or organizations assign to specific values. Traditionally, it relies on tools like self-report questionnaires [80, 53], behavioral observation [51, 54, 23], and experimental techniques [71, 87, 57]. However, these tools are hindered by response biases, resource demands, inaccuracies in capturing authentic behaviors, and inability to handle historical or subjective data [60, 10, 8]. Recently proposed Generative Psychometrics for Values (GPV) [92] leverage LLMs to dynamically parse unstructured texts into perceptions and measure values therein. GPV is shown to be more scalable, reliable, and flexible than traditional tools in measuring LLM values, and is utilized for non-reactive value measurement in this work. For more details, please refer to § A.3 or the original literature [92].
@@
Isolated values are structured into a value system for a holistic understanding of their relationships and implications [76]. Theory-driven value systems like Schwartz’s theory of basic human values [75] and functional theory of values [27] are rooted in theoretical hypotheses and preconceptions of psychologists. These systems are established prior to data collection and analysis, and are later validated through empirical studies. However, theory-driven approaches are constrained by the subjectivity of the theorists, limited coverage, and an inability to adapt to evolving values or specific contexts [60, 61].
@@ 2.2 Psycho-Lexical Approach
In contrast to the theory-driven approaches, the psycho-lexical approach organizes psychological constructs such as values, personality traits, and social attitudes [1, 16, 42, 72], under a data-driven paradigm. It operates on the fundamental principle that language naturally evolves to capture salient and socially relevant individual differences [18]. The pioneering attempts date back to the early 20th century [2, 24], and the paradigm has then been developed and extended to different psychological constructs for decades [14, 40, 1, 60, 52, 26].
@@ 2.3 From Human Values to LLM Values
The rise of LLMs also brings the critical need to evaluate, understand, and align their values. Extensive works focus on evaluating the values of LLMs using traditional self-report tools [48, 49, 32, 68, 69], static customized inventories [65, 55, 36], dynamically generated inventories [37], or directly from the model’s free-form outputs [92, 90, 89, 91]. Other works attempt to understand the LLM values in aspects like the consistency of shown values [67, 56, 66], the ability to reason about human values [25, 82, 39, 84], and the ability to express or role-play human values [38, 94, 41, 93, 46, 33]. Another line of research aligns LLM values with human values, setting the alignment goal as specific risk metrics [50, 30], human demonstrations [22, 43, 86], implicit preference modeling [59, 62, 96], or intrinsic values [4, 89, 6]. Among different alignment goals, intrinsic values, as transituational decision-guiding principles, demonstrate unique advantages [88]. However, prior works are mostly based on Schwartz’s value system, which is established and validated for humans and may not necessarily capture the unique LLM psychology. Therefore, it is essential to construct a value system for LLMs grounded in psychological theories and principles, based on which we can more effectively evaluate, understand, and align the values of LLMs.
One preliminary attempt [9] of constructing an LLM value system faces several limitations that undermine its psychological grounding: 1) the value lexicon collection is prone to response bias and lacks comprehensive coverage; 2) the derivation of value correlations is not theoretically grounded and lacks empirical validity; and 3) the resulting value system has not been evaluated in terms of its reliability, validity, or utility. Appendix C presents further discussion and experimental evidence.
@@ 3 Generative Psycho-Lexical Approach for Constructing LLM Value System
In line with the established psycho-lexical approach [18], we begin by comprehensively identifying atomic values that LLMs may endorse. We draw upon a diverse, value-laden corpus of LLM outputs $\mathcal{C}$ from ValueBench [65], GPV [92], BeaverTails [35], and ValueLex [9]. These corpora provide a broad spectrum of value-laden LLM generations. Detailed statistics and further explanations on these datasets are provided in § A.1.
@@
Then, we extract perceptions $P$ from the collected corpus using the Perception Parser $M_{P}$. Perceptions are atomic value-rich expressions that reflect values behind free-form LLM outputs and are akin to stimuli in traditional psychometric tools [92]. These perceptions are then mapped to underlying values $V$ through the Value Generator $M_{G}$, with their frequency of occurrence $Freq$ recorded. We instantiate $M_{P}$ following GPV [92] and $M_{G}$ using Kaleido model [82]. In § B.1, we confirm the comprehensive coverage of the generated values.
@@
To ensure the value lexicons are concise and representative, we filter out duplicates using Rouge scores and embedding similarity following the validated setup in [82]. When two values exhibit significant semantic overlap, we retain the one more frequently observed, as higher frequency signifies greater priority [60]. This process results in value lexicons $V$ that minimize redundancy while preserving semantic coverage.
@@
Structuring the LLM value system requires identifying correlations between values derived from LLM value measurements. To this end, we adopt GPV [92], instantiated with $M_{P}$ and $M_{E}$, to measure the value orientations $\mathbf{X}_{M}$ of LLM subjects $M$ on values $V$. GPV dynamically extracts psychometric stimuli from LLM outputs and reveals value orientations therein. We measure a set of 693 LLM subjects, consisting of 33 LLMs paired with 21 profiling prompts (§ A.2). The involved LLMs are instruction-tuned models, as they are more relevant for real-world, public-facing applications. The value system $\mathcal{VS}=(V,V_{H},\mathbf{\Lambda})$ is then constructed through Principal Component Analysis (PCA) [60], which uncovers the underlying value factors $V_{H}$ and hierarchical relationships $\mathbf{\Lambda}$ (loadings of $V$ on $V_{H}$).
@@ 4.1 Are LLMs Capable for GPLA?
Perceptions are value-laden expressions dynamically extracted from LLM outputs. They resemble the stimuli in traditional psychometric tools [92]. The Perception Parser $M_{P}$ extracts such perceptions from the LLM outputs. As validated by Ye et al. [92], $M_{P}$ can extract high-quality perceptions suitable for subsequent analysis. In addition, related research confirms that LLMs demonstrate abilities beyond human golden standards in value-related generation tasks [65, 82, 97].
The Value Generator $M_{G}$ identifies values from the extracted perceptions. As validated by Sorensen et al. [82], $M_{G}$, instantiated with the Kaleido model, can generate values that are both comprehensive and high-quality, with 91% of the generations marked as good by all three annotators, and missing values detected 0.35% of the time.
The Value Evaluator $M_{E}$ measures the value orientations of LLMs given the extracted perceptions. When instantiated with ValueLlama [92], $M_{E}$ can approximate the psychologists’ judgments on held-out values with about 90% accuracy.
@@ 4.2 Advantages of GPLA over Prior Psycho-Lexical Approaches
Here we discuss the main advantages of GPLA over the traditional psycho-lexical approach in constructing value systems. We defer the comparison with ValueLex [9], a recent study that constructs LLM value systems, to Appendix C.
@@
  • Prioritization Criteria. The scalable collection of value lexicons by LLMs enables the concurrent collection of their occurrence frequencies. Since these frequencies reflect the salience of values [60], they can serve as criteria for prioritizing values during the filtering process.
@@
Traditional methods of value measurement typically rely on self-reports or peer ratings, which are prone to response biases, resource-intensive, and often fail to capture authentic behaviors or account for historical and subjective data [60]. In contrast, GPLA utilizes non-reactive and LLM-driven value measurement, which is proven to be more scalable and reliable [92].
@@ 5 Benchmarking LLM Value Systems
The power of a value system lies in its validity and utility [76]. This work introduces, for the first time, three benchmarking tasks for LLM value systems, rooted in both psychological theories and AI priorities: (1) Confirmatory Factor Analysis, which evaluates their structure validity given LLM value measurements; (2) LLM Safety Prediction, which assesses their predictive validity for LLM safety; and (3) LLM Value Alignment, which examines their representation power in aligning LLMs with human values.
@@ 5.1 Confirmatory Factor Analysis for Evaluating Structure Validity
Confirmatory Factor Analysis (CFA) is a statistical technique used to test whether a set of observed variables (atomic values) accurately represents and loads onto a smaller number of underlying latent factors (high-level values) [79]. A well-fitted CFA model confirms the theoretical structure of a value system, by showing that the observed measurements of atomic values are consistent with their expected relationships to the high-level latent value factors.
Mathematically, the CFA model is represented as:
@@
where $\mathbf{X}\in\mathbb{R}^{n\times|V|}$ is the observed data matrix (measuring atomic values); $\mathbf{F}\in\mathbb{R}^{n\times|V_{H}|}$ is the matrix of latent factors (aggregated measurements of high-level value factors); $\mathbf{\Lambda}\in\mathbb{R}^{|V|\times|V_{H}|}$ is the matrix of factor loadings (the contribution of each atomic value to each latent factor); and $\mathbf{\epsilon}\in\mathbb{R}^{n\times|V|}$ is the matrix of error terms.
On held-out LLM value measurement data, this task evaluates the model fit using metrics such as the Root Mean Square Error of Approximation (RMSEA). Better model fit indicates higher structure validity of the value system.

### v2

@@ Abstract
[…]
[table]capposition=bottom
Generative Psycho-Lexical Approach for Constructing Value Systems in Large Language Models
Anonymous ACL submission
@@ 1 Introduction
Personal values are broad, fundamental motivations behind individual and collective actions. They serve as guiding principles that shape perceptions, cognition, and behaviors across situations and over time. As Large Language Models (LLMs) continue to permeate and transform human society, researchers have increasingly sought to evaluate, understand, and align LLM values (ye2025gpv; ren2024valuebench; rao2024normad; meadows2024localvaluebench; kovavc2024stick; jiang2024raising; rozen2024llms; yao2024clave; zhang2023measuring; sorensen2024value; moore2024large; demszky2023using). Rather than relying solely on specific risk metrics, implicit preference modeling, or atomic ethical principles, framing these studies through the lens of intrinsic values schwartz2012overview offers a more comprehensive, adaptable, and transparent approach yao2023value_fulcra; biedma2024beyond; yao2023alignment_goals.
@@
However, value systems used in existing research—primarily Schwartz’s values—are designed for humans. A psychologically informed value system specifically tailored for LLMs remains largely unexplored, obscuring a structured and holistic perspective on their values. A preliminary attempt at constructing LLM value systems biedma2024beyond has several limitations that undermine its psychological grounding; we present related discussion and experimental evidence in § 2.3 and Appendix C.
@@
To address this gap, we introduce a generative psycho-lexical approach (GPLA), a novel methodology that leverages LLMs to construct a value system grounded in psychological principles. This approach is based on the psycho-lexical hypothesis, which suggests that all salient values are captured in language. GPLA adopts an agentic framework and involves five sequential steps: 1) extracting perceptions from the corpus, 2) identifying values behind perceptions, 3) filtering values, 4) non-reactive value measurements of test subjects, and 5) structuring the value system. In contrast to traditional methods that manually compile value lexicons and design questionnaires to collect self-report data aavik2002structure; crectu2012psycho; Klages1929-KLATSO-5; saucier2000isms, GPLA automates the entire process and features two key advantages.
@@
First, GPLA supports agent-specific construction. Rather than relying on value lexicons derived from human-compiled dictionaries, it utilizes LLMs to identify values within free-form agent-generated text. This principled approach facilitates the agent- and context-specific collection of value lexicons and ensures better coverage (§ B.1). Second, GPLA enables nonreactive value measurement. Traditional methods that rely on forced-choice self-report are demonstrated to be unreliable and biased for humans ponizovskiy2020development; bardi2008new and AI ye2025gpv; rottger2024political; dominguez2023questioning. Instead, GPLA measures values through analyzing unstructured responses to arbitrary prompts, which mitigates response biases and enables scalable and self-adaptive value measurement ye2025gpv.
@@
In applying GPLA to 33 LLMs, each steered with 21 different value-anchoring prompts rozen2024llms, we develop a five-factor value system encompassing Social Responsibility, Risk-taking, Rule-following, Self-competence, and Rationality. We demonstrate the system’s reliability and explore its theoretical and practical implications. To benchmark LLM value systems, we propose three tasks based on psychometric principles devon2007psychometric and AI priorities ji2023ai_alignment; anwar2024foundational: Confirmatory Factor Analysis for assessing structural validity, LLM Safety Prediction for examining predictive validity, and LLM Value Alignment for evaluating representational power. These tasks confirm the validity and utility of our system in evaluating, understanding, and aligning the intrinsic values of LLMs, when compared to canonical Schwartz’s values schwartz2012overview.
@@ 2.1 Values, Value Measurement, and Theory-Driven Value Systems
Values are broad, fundamental beliefs or principles that guide individuals’ behaviors, cognition, and decisions schwartz1987toward. They represent what people consider important in life and serve as motivational forces when evaluating actions, people, and events schwartz2012overview. Unlike other psychological constructs such as personalities, values are stable, motivational, and transituational sagiv2022personal. These characteristics make values uniquely powerful for understanding human behaviors bardi2003values, evaluating individual and cultural differences schwartz2006theory, and resolving conflicts rooted in clashing priorities sen1986social.
@@
Values are measurable constructs schwartz2001extending. Value measurement seeks a quantitative assessment of the importance individuals or organizations assign to specific values. Traditionally, it relies on tools like self-report questionnaires schwartz2001extending; maio2010mental, behavioral observation lonnqvist2013personal; maio2009changing; fischer2011whence, and experimental techniques sagiv2011compete; yamagishi2013behavioral; murphy2014social. However, these tools are hindered by response biases, resource demands, inaccuracies in capturing authentic behaviors, and inability to handle historical or subjective data ponizovskiy2020development; boyd2015values; bardi2008new. Recently proposed Generative Psychometrics for Values (GPV) ye2025gpv leverage LLMs to dynamically parse unstructured texts into perceptions and measure values therein. GPV is shown to be more scalable, reliable, and flexible than traditional tools in measuring LLM values, and is utilized for non-reactive value measurement in this work.
@@
Isolated values are structured into a value system for a holistic understanding of their relationships and implications schwartz2012overview. Theory-driven value systems like Schwartz’s theory of basic human values schwartz1992universals and functional theory of values gouveia2014functional are rooted in theoretical hypotheses and preconceptions of psychologists. These systems are established prior to data collection and analysis, and are later validated through empirical studies. However, theory-driven approaches are constrained by the subjectivity of the theorists, limited coverage, and an inability to adapt to evolving values or specific contexts ponizovskiy2020development; raad2017reply.
@@ 2.2 Psycho-Lexical Approach
In contrast to the theory-driven approaches, the psycho-lexical approach organizes psychological constructs such as values, personality traits, and social attitudes aavik2002structure; crectu2012psycho; Klages1929-KLATSO-5; saucier2000isms, under a data-driven paradigm. It operates on the fundamental principle that language naturally evolves to capture salient and socially relevant individual differences de2016values. The pioneering attempts date back to the early 20th century allport1936trait; galton1950character, and the paradigm has then been developed and extended to different psychological constructs for decades cattell1957personality; john1988lexical; aavik2002structure; ponizovskiy2020development; mai2023exploring; garrashi2024personality.
@@ 2.3 From Human Values to LLM Values
The rise of LLMs also brings the critical need to evaluate, understand, and align their values. Extensive works focus on evaluating the values of LLMs using traditional self-report tools li2022gpt; li2024evaluatingpsychologicalsafetylarge; huang2024humanity; safdari2023-personality-traits-in-llm; safdari2023personality, static customized inventories ren2024valuebench; meadows2024localvaluebench; jiang2024evaluating, dynamically generated inventories jiang2024raising, or directly from the model’s free-form outputs ye2025gpv; yao2024clave; yao2023value_fulcra; yao2025leaderboard. Other works attempt to understand the LLM values in aspects like the consistency of shown values rozen2024llms; moore2024large; rottger2024political, the ability to reason about human values ganesan2023systematic; sorensen2024value; jiang2024can; strachan2024testing, and the ability to express or role-play human values jiang-etal-2024-personallm; zhang2023measuring; kang2024causal; zhang2025extrapolating; li2024quantifying; huang2024social. Another line of research aligns LLM values with human values, setting the alignment goal as specific risk metrics lin2024towards; gunjal2024detecting, human demonstrations dubois2024alpacafarm; kopf2024openassistant; alpaca, implicit preference modeling ouyang2022training; rafailov2024dpo; zhong2024panacea, or intrinsic values bai2022hh; yao2023value_fulcra; bai2022constitutional. Among different alignment goals, intrinsic values, as transituational decision-guiding principles, demonstrate unique advantages yao2023alignment_goals. However, prior works are mostly based on Schwartz’s value system, which is established and validated for humans and may not necessarily capture the unique LLM psychology. Therefore, it is essential to construct a value system for LLMs grounded in psychological theories and principles, based on which we can more effectively evaluate, understand, and align the values of LLMs.
One preliminary attempt biedma2024beyond of constructing an LLM value system faces several limitations that undermine its psychological grounding: 1) the value lexicon collection is prone to response bias and lacks comprehensive coverage; 2) the derivation of value correlations is not theoretically grounded and lacks empirical validity; and 3) the resulting value system has not been evaluated in terms of its reliability, validity, or utility. Appendix C presents further discussion and experimental evidence.
@@ 3 Generative Psycho-Lexical Approach for Constructing LLM Value System
In line with the established psycho-lexical approach de2016values, we begin by comprehensively identifying atomic values that LLMs may endorse. We draw upon a diverse, value-laden corpus of LLM outputs $\mathcal{C}$ from ValueBench ren2024valuebench, GPV ye2025gpv, BeaverTails ji2024beavertails, and ValueLex biedma2024beyond. These corpora provide a broad spectrum of value-laden LLM generations. Detailed statistics and further explanations on these datasets are provided in § A.1.
@@
Then, we extract perceptions $P$ from the collected corpus using the Perception Parser $M_{P}$. Perceptions are atomic value-rich expressions that reflect values behind free-form LLM outputs and are akin to stimuli in traditional psychometric tools ye2025gpv. These perceptions are then mapped to underlying values $V$ through the Value Generator $M_{G}$, with their frequency of occurrence $Freq$ recorded. We instantiate $M_{P}$ following GPV ye2025gpv and $M_{G}$ using Kaleido model sorensen2024value. In § B.1, we confirm the comprehensive coverage of the generated values.
@@
To ensure the value lexicons are concise and representative, we filter out duplicates using Rouge scores and embedding similarity following the validated setup in sorensen2024value. When two values exhibit significant semantic overlap, we retain the one more frequently observed, as higher frequency signifies greater priority ponizovskiy2020development. This process results in value lexicons $V$ that minimize redundancy while preserving semantic coverage.
@@
Structuring the LLM value system requires identifying correlations between values derived from LLM value measurements. To this end, we adopt GPV ye2025gpv, instantiated with $M_{P}$ and $M_{E}$, to measure the value orientations $\mathbf{X}_{M}$ of LLM subjects $M$ on values $V$. GPV dynamically extracts psychometric stimuli from LLM outputs and reveals value orientations therein. We measure a set of 693 LLM subjects, consisting of 33 LLMs paired with 21 profiling prompts (§ A.2). The involved LLMs are instruction-tuned models, as they are more relevant for real-world, public-facing applications. The value system $\mathcal{VS}=(V,V_{H},\mathbf{\Lambda})$ is then constructed through Principal Component Analysis (PCA) ponizovskiy2020development, which uncovers the underlying value factors $V_{H}$ and hierarchical relationships $\mathbf{\Lambda}$ (loadings of $V$ on $V_{H}$).
@@ 4.1 Are LLMs Capable for GPLA?
Perceptions are value-laden expressions dynamically extracted from LLM outputs. They resemble the stimuli in traditional psychometric tools ye2025gpv. The Perception Parser $M_{P}$ extracts such perceptions from the LLM outputs. As validated by ye2025gpv, $M_{P}$ can extract high-quality perceptions suitable for subsequent analysis. In addition, related research confirms that LLMs demonstrate abilities beyond human golden standards in value-related generation tasks ren2024valuebench; sorensen2024value; ziems2024can.
The Value Generator $M_{G}$ identifies values from the extracted perceptions. As validated by sorensen2024value, $M_{G}$, instantiated with the Kaleido model, can generate values that are both comprehensive and high-quality, with 91% of the generations marked as good by all three annotators, and missing values detected 0.35% of the time.
The Value Evaluator $M_{E}$ measures the value orientations of LLMs given the extracted perceptions. When instantiated with ValueLlama ye2025gpv, $M_{E}$ can approximate the psychologists’ judgments on held-out values with about 90% accuracy.
@@ 4.2 Advantages of GPLA over Prior Psycho-Lexical Approaches
Here we discuss the main advantages of GPLA over the traditional psycho-lexical approach in constructing value systems. We defer the comparison with ValueLex biedma2024beyond, a recent study that constructs LLM value systems, to Appendix C.
@@
  • Prioritization Criteria. The scalable collection of value lexicons by LLMs enables the concurrent collection of their occurrence frequencies. Since these frequencies reflect the salience of values ponizovskiy2020development, they can serve as criteria for prioritizing values during the filtering process.
@@
Traditional methods of value measurement typically rely on self-reports or peer ratings, which are prone to response biases, resource-intensive, and often fail to capture authentic behaviors or account for historical and subjective data ponizovskiy2020development. In contrast, GPLA utilizes non-reactive and LLM-driven value measurement, which is proven to be more scalable and reliable ye2025gpv.
@@ 5 Benchmarking LLM Value Systems
The power of a value system lies in its validity and utility schwartz2012overview. This work introduces, for the first time, three benchmarking tasks for LLM value systems, rooted in both psychological theories and AI priorities: (1) Confirmatory Factor Analysis, which evaluates their structure validity given LLM value measurements; (2) LLM Safety Prediction, which assesses their predictive validity for LLM safety; and (3) LLM Value Alignment, which examines their representation power in aligning LLMs with human values.
@@ 5.1 Confirmatory Factor Analysis for Evaluating Structure Validity
Confirmatory Factor Analysis (CFA) is a statistical technique used to test whether a set of observed variables (atomic values) accurately represents and loads onto a smaller number of underlying latent factors (high-level values) schwartz2004cfa. Mathematically, the CFA model is represented as:
@@
[…]
This task evaluates the model fit on held-out value measurement data. Better model fit indicates higher structure validity of the system.